a gas container contains 35.0 l at 150 bar determine the volume in l this gas will occupy if it were allowed to expand at atmospheric temperature room temperature


1
Expert's answer
2021-12-14T16:29:01-0500

V1 = 35L

P1= 150bar

At standard condition, pressure, P2 = 1.01325 bar

V2 =?

Using Boyle's law, P1V1=P2V2

150*35 = 1.01325*V2

V2 = (150*35)/1.01325 = 5181.35L